# revtex4-2Repair the float

### v2

@@ Abstract
We present an effective field theory for a unified description of transport in normal and superconducting metals in the presence of generic spin-orbit coupling (SOC). The structure of the quantum kinetic theory in the diffusive regime is determined by a set of fundamental constraints – charge conjugation symmetry, the causality principle, and the crystal symmetry of a material. These symmetries uniquely fix the action of the Keldysh non-linear $\sigma$ model (NLSM), which at the saddle point yields the quantum kinetic Usadel-type equation. Our phenomenological approach is reminiscent of the Ginzburg-Landau theory, but is valid for superconductors in the whole temperature range, describes the diffusive transport in the normal state, and naturally captures the effects of superconducting fluctuations. As an application, we derive the NLSM and quantum transport equations which include all effects of spin-orbit coupling, allowed by the crystal symmetry, for example, the spin Hall, spin current swapping or spin-galvanic effects. Our approach can be extended to systems with broken time reversal symmetry, as well as to the description of hybrid interfaces, where the spin-charge interconversion can be enhanced due to strong interfacial SOC.
Electronic transport in materials with spin-orbit coupling (SOC) has attracted a lot of interest in recent decades, particularly due to the possibility of spin-charge interconversion [1, 2]. This interest has extended to systems where superconductivity coexists with spin-orbit coupling, motivated by intensive research in the fields of topological superconductivity [3], two-dimensional superconductors and van der Waals heterostructures [4], and non-centrosymmetric superconductors [5]. The interplay between superconductivity and spin-orbit coupling underlies the physics of superconducting magnetoelectric and spin diode effects [6, 7].
@@
The central kinetic equation in diffusive superconductors is the Usadel equation [23]. In systems without spin-orbit coupling, it can be derived directly from the Gorkov equations following a standard procedure [24]. However, mixing of charge and spin degrees of freedom in the presence of small, but finite spin-orbit coupling (SOC) prevents a straightforward application of such a standard scheme. Recently, it has been realized [15, 16] that in this case, the Usadel equation can be conveniently derived from the non-linear sigma model (NLSM) [25, 26, 27], as a saddle point of effective action for disordered conductors. In Refs. [15, 16] the NLSM was derived for the special cases of extrinsic and Rashba-like intrinsic SOC, following the standard microscopic procedure of averaging over Gaussian disorder (see, for example, Refs. [28, 29, 30] for different versions of the NLSM). Apparently, such an approach can only be effective for simplified models, while for realistic materials it becomes prohibitively tedious because of their complex spin-dependent electronic structure and the necessity to integrate various mechanisms of SOC.
Here we propose another approach that, similarly to the Ginzburg-Landau (GL) theory, relies solely on symmetry arguments. This allows us to incorporate effects of spin-orbit coupling into the NLSM and formulate a phenomenological quantum kinetic theory. Our theory captures all symmetry-allowed transport effects associated with spin, charge, and their coupling, such as spin-Hall and spin-galvanic effects. In contrast to GL theory, this phenomenology is valid for arbitrary temperatures, away from equilibrium, and covers the normal state. In the following, we identify the basic symmetries and explain how to construct the generalized NLSM. Then, the saddle point condition leads us to our main objective: obtaining the general quantum kinetic equation for disordered conductors with spin-orbit coupling, which reveals all physical effects associated with the different terms of the action.
@@
where the Pauli matrices $\sigma_{i}$ and $\tau_{i}$ span the spin and the Nambu spaces, respectively, $\hat{\omega}_{tt^{\prime}}=\partial_{t}\delta(t-t^{\prime})$ is the time derivative operator, $D$ is the diffusion coefficient, and $\hat{\Delta}=\tau_{2}|\Delta|e^{i\tau_{3}\varphi}$ is the superconducting order parameter matrix. The second line in Eq. (1) describes the effects of extrinsic SOC: the spin Hall effect, spin current swapping, and the spin relaxation, parameterized by the Hall angle $\theta$, the swapping coefficient $\varkappa$, and the spin relaxation rate $\Gamma$, respectively. The external electromagnetic potential is introduced through the replacement $\partial_{\mu}\mapsto\partial_{\mu}-i[A_{\mu}\tau_{3},.]$, and the Zeeman or exchange field is denoted by $\mathbf{h}$.
@@
The first general symmetry of the NLSM is related to the redundancy of the Hilbert space in the Nambu description of superconductors, which involves both the full set of electronic states and their time-reversal conjugated counterparts. In non-superconducting systems, this technical doubling of the number of degrees of freedom is a convenient way to include fluctuations in the Cooperon channel. Because of the above redundancy, not all components of the matrix $Q$-field in NLSM are independent, but are constrained to satisfy the charge conjugation symmetry [25, 29, 30]. In the Keldysh contour representation, this fundamental symmetry condition reads $Q=C{Q}^{T}C^{-1}$, where $C=-{i}\sigma_{2}\tau_{1}$ is the charge conjugation operator, and the transposition operation involves both the usual matrix transposition and the interchanging of the contour time arguments  [^2] It is worth noting that the charge conjugation is an exact and universal symmetry of the Bogolyubov-de Gennes equations, which is the consequence of the redundancy of the Nambu formalism [50, 51, 52].. In the Keldysh matrix representation, after the standard Keldysh rotation [30] the charge conjugation symmetry of the 8$\times$8 matrix $Q$-field takes the form (see SM Sec. A for details),
@@
To illustrate this procedure, we derive contributions to the NLSM which may appear in the presence of SOC. More details are presented in SM Sec. C. Physically, SOC mixes the spin and orbital/translation degrees of freedom, resulting in the simultaneous presence of $\sigma$-matrices and spatial gradients in the NLSM. We construct the nontrivial contributions with the lowest orders on both Pauli matrices and gradients and discuss their main physical effects. Since SOC preserves the time-reversal, we only consider terms invariant under the operation $Q\mapsto\tau_{3}\sigma_{y}Q^{T}\sigma_{y}\tau_{3}$.
Second order – spin precession and relaxation. In the second order, there are only three primitives consistent with the condition $Q^{2}=1$ and Eq. (2), namely, ${\rm Tr}(D_{kj}\partial_{k}Q\partial_{j}Q)$, ${\rm Tr}(\Gamma_{kj}\sigma_{k}Q\sigma_{j}Q)$, and ${\rm Tr}(\tilde{\alpha}_{kj}\sigma_{k}Q\partial_{j}Q)$. The chronological symmetry of Eq. (3) requires that $D_{kj}$ and $\Gamma_{kj}$ are real symmetric tensors, while $\tilde{\alpha}_{kj}=i\alpha_{kj}$ is an imaginary pseudotensor. $D_{kj}$ and $\Gamma_{kj}$ correspond, respectively, to the diffusion coefficient and the spin relaxation rate, which can be anisotropic, depending on the crystal symmetry. A second rank pseudotensor $\alpha_{kj}$ is allowed only in gyrotropic materials [37] and the corresponding term is responsible for the spin precession induced by SOC. By parameterizing $\alpha_{jk}$ as follows, $\alpha_{jk}=4\mathcal{A}_{jl}D_{lk}$ the second order part of the NLSM can be cast in a compact form of SU(2) covariant diffusion with an additional spin relaxation,
$iS_{2}=\frac{\pi\nu}{8}{\rm Tr}\Big{(}-D_{kj}\tilde{\nabla}_{k}Q\tilde{\nabla}_{j}Q+\frac{1}{4}\tilde{\Gamma}_{kj}\sigma_{k}Q\sigma_{j}Q\Big{)}$ (4)
where ~$\tilde{\nabla}_{k}=\partial_{k}-i[\hat{\mathcal{A}}_{k},\cdot]$ is the SU(2) covariant derivative with an effective gauge field $\hat{\mathcal{A}}_{k}=\mathcal{A}_{jk}\sigma_{j}$, and $\tilde{\Gamma}_{kj}=\Gamma_{kj}-4D_{ab}\mathcal{A}_{ka}\mathcal{A}_{jb}$  [^4] We note that stability of the system require $D_{kj}$ and $\Gamma_{kj}$ to be positive definite.. The SU(2) covariant diffusion has been introduced microscopically [11, 39] for a special class of $\mathbf{k}\cdot\mathbf{p}$ models with a Rashba-like linear SOC. Now we see that it is generic for gyrotropic materials. In realistic systems, the SU(2) symmetry of simple models is of course broken, but at the leading order this leads only to extra spin relaxation.
Third order – spin-charge coupling related terms. There are two types of the third order terms:
@@
$iS_{3}^{(2)}=\frac{\pi\nu}{16}{\rm Tr}\Big{(}i\varepsilon_{ijl}\gamma_{lk}\sigma_{i}Q\sigma_{j}Q\partial_{k}Q+\beta_{ijk}\sigma_{i}Q\sigma_{j}\partial_{k}Q\Big{)}$ (6)
@@
$\nabla_{k}\mathcal{J}_{k}+[\tau_{3}(\hat{\omega}+i\mathbf{h}{\bm{\sigma}})+\hat{\Delta},g]=\mathcal{T}-\frac{1}{8}\Gamma_{jk}[\sigma_{j}g\sigma_{k},g]$ (7)
@@
$\mathcal{J}_{k}$ $=-Dg\nabla_{k}g+\frac{i}{16}\varepsilon_{ijl}\gamma_{lk}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}\,,$ (8)
$\mathcal{T}$ $=-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\,.$ (9)
By construction, these equations are consistent with the normalization condition $g^{2}=1$. Importantly, anticommutators of $g$ with spin matrices in Eqs. (8)-(9) induce coupling between singlet and triplet components of $g$. This indicates that the pseudotensor $\gamma_{ij}$ is related to magnetoelectric phenomena. The precise physics encoded in $\gamma_{ij}$ can be understood by analyzing Eqs. (8)-(9) in the linear regime.
@@
The second term in Eq. (10) describes a spin independent contribution to the matrix current proportional to the triplet (spin dependent) part of $\delta g$. Apparently, this corresponds to the spin-galvanic effect (SGE), also called the inverse Edelstein effect [2, 40, 41, 42, 43, 44, 45, 46, 47]. Indeed, the physical current is obtained from the Keldysh component of the matrix current $j_{k}=\frac{\pi\nu}{2}{\rm Tr}\{\tau_{3}\mathcal{J}_{k}^{K}(t,t)\}$, which gives a linear relation between the charge current and the excess spin $\delta\mathbf{S}=\frac{\pi\nu}{2}{\rm Tr}\{\tau_{3}\mathbf{g}_{t}^{K}(t,t)\}$,
$j_{k}^{\rm SGE}=\gamma_{ik}\delta S_{i}$ (12)
@@
In a vortex, the phase of the pair amplitude only depends on the angle, while their magnitude depends on the radial coordinate $r$. As elaborated in SM Sec. D, in Eq. (11) the phase derivative enters as an effective exchange field $\gamma_{kj}\hat{\phi}_{k}[\sigma_{j}\tau_{3},g_{s}]$, where $\bm{\hat{\phi}}$ denotes the unit vector in the tangential direction, while the derivative with respect to the radial coordinate does not break time-reversal symmetry and hence does not contribute to the spin texture. Thus, the spin around the vortex points along $\gamma_{kj}\hat{\phi}_{k}$. The inset of Fig. 1 sketches the textures for different point group symmetries. The dependence of the current density and induced spin density as a function of the radial coordinate, has to be computed numerically by solving the Usadel equation, see SM Sec. D for details. Fig. 1 shows the result for temperatures well below the critical temperature. It is clear that the relation between the two is non-local, since $S(r)$ is maximized further away from the core than the maximum current density. This non-locality is reduced by increasing the temperature.
@@
In the limit of high temperature, $T\lesssim T_{c}$ the GL functional can be straightforwardly derived. In a gyrotropic system, the term $q_{j}\gamma_{ij}$ acts as an effective exchange field, which combines with $\mathbf{h}_{i}$ (where $q_{k}$ represents the Fourier transform of the derivatives $i\partial_{k}$; see SM Sec. E). Consequently, the GL free energy includes a term linear in derivatives known as the Lifshitz invariant, denoted as $F_{L}=\chi_{ij}h_{j}\gamma_{ik}\partial_{k}\varphi$, where $\chi_{ij}$ is the magnetic response function to a Zeeman field, inducing an excess spin, $\delta S_{i}=\chi_{ij}h_{j}$. This establishes a connection of the Lifshitz invariant to the universal spin-galvanic coefficient $\gamma_{ij}$ appearing in Eq. (12) and in the action Eq. (6). The full list of all Lifshitz invariants for all gyrotropic point groups has been presented, for example, in Ref. [49]
In conclusion, we have presented a method for deriving the action of the Keldysh NLSM exclusively relying on the principles of charge conjugation, chronological and crystal symmetries. Applying this method, we have successfully derived the quantum kinetic equation, commonly known as the Usadel equation, for metallic systems featuring generic SOC. Through this derivation, we have identified the distinct terms and symmetries governing spin precession, relaxation, and spin-charge coupling. As a byproduct, the provided NLSM can be used to investigate superconducting fluctuations or localization effects in materials with arbitrary types of SOC. Our approach, while akin in spirit to the GL theory, offers an effective field theory valid in the whole temperature range. Moreover, it establishes a novel framework for deriving quantum kinetic theories applicable to diverse systems.

### v3

@@ Abstract
We present a unified description of transport in normal and superconducting metals in the presence of generic spin-orbit coupling (SOC). The structure of the quantum kinetic theory in the diffusive regime is determined by a set of fundamental constraints – charge conjugation symmetry, the causality principle, and the crystal symmetry of a material. These symmetries uniquely fix the action of the Keldysh nonlinear $\sigma$ model (NLSM), which at the saddle point yields the quantum kinetic Usadel-type equation, the equation that describes the main transport features of a system. Our phenomenological approach is reminiscent of the Ginzburg-Landau theory, but is valid for superconductors in the whole temperature range, describes the diffusive transport in the normal state, and naturally captures the effects of superconducting fluctuations. As an application, we derive the NLSM and the corresponding quantum transport equations which include all effects of spin-orbit coupling, allowed by the crystal symmetry, for example, the spin Hall, spin current swapping or spin-galvanic effects. Our approach can be extended to derive transport equations in systems with broken time reversal symmetry, as well as to the description of hybrid interfaces, where the spin-charge interconversion can be enhanced due to strong interfacial SOC.
Electronic transport in materials with spin-orbit coupling (SOC) has attracted a lot of interest in recent decades, particularly due to the possibility of spin-charge interconversion [1, 2]. This interest has extended to systems where superconductivity coexists with spin-orbit coupling, motivated by intensive research in the fields of topological superconductivity [3], two-dimensional superconductors and van der Waals heterostructures [4], and noncentrosymmetric superconductors [5]. The interplay between superconductivity and spin-orbit coupling underlies the physics of superconducting magnetoelectric and spin diode effects [6, 7].
@@
The central kinetic equation in diffusive superconductors is the Usadel equation [23]. In systems without spin-orbit coupling, it can be derived directly from the Gorkov equations following a standard procedure [24, 25]. However, mixing of charge and spin degrees of freedom in the presence of small, but finite spin-orbit coupling (SOC) prevents a straightforward application of such a standard scheme. Recently, it has been realized [15, 16] that in this case, the Usadel equation can be conveniently derived from the nonlinear sigma model (NLSM) [26, 27, 28], as a saddle point of effective action for disordered conductors. In Refs. [15, 16] the NLSM was derived for the special cases of extrinsic and Rashba-like intrinsic SOC, following the standard microscopic procedure of averaging over Gaussian disorder (see, for example, Refs. [29, 30, 31, 32] for different versions of the NLSM). Apparently, such an approach can only be effective for simplified models, while for realistic materials it becomes prohibitively tedious because of their complex spin-dependent electronic structure and the necessity to integrate various mechanisms of SOC.
Here we propose another approach to derive the corresponding quantum kinetic equations via the NLSM, that, similarly to the Ginzburg-Landau (GL) theory, relies solely on symmetry arguments. This allows us to incorporate effects of spin-orbit coupling into the NLSM and formulate a phenomenological quantum kinetic theory. Our theory captures all symmetry-allowed transport effects associated with spin, charge, and their coupling, such as spin-Hall and spin-galvanic effects. In contrast to GL theory, this phenomenology is valid for arbitrary temperatures, away from equilibrium, and covers the normal state. In the following, we identify the basic symmetries and explain how to construct the generalized NLSM. Then, the saddle point condition leads us to our main objective: obtaining the general quantum kinetic equation for disordered conductors with spin-orbit coupling, which reveals all physical effects associated with the different terms of the action.
The obtained quantum kinetic equation can be used for a broad scope of materials. In the specific cases of an SU(2) covariant Rashba type spin-orbit coupling or exclusively isotropic extrinsic SOC our model recovers the equations found in [16] and [15] respectively. However, the presented equations have a far more general scope and can be used as well for materials with different types of spin-orbit coupling and the combination of several sources of SOC.
@@
where the Pauli matrices $\sigma_{i}$ and $\tau_{i}$ span the spin and the Nambu spaces, respectively, $\hat{\omega}_{tt^{\prime}}=\partial_{t}\delta(t-t^{\prime})$ is the time derivative operator, $D$ is the diffusion coefficient, and $\hat{\Delta}=\tau_{2}|\Delta|e^{i\tau_{3}\varphi}$ is the superconducting order parameter matrix which pairs time-reversed partners [41]. The second line in Eq. (1) describes the effects of extrinsic SOC: the spin Hall effect, which connects mutually orthogonal charge and spin currents, the spin current swapping [42], which mixes spin currents with ”swapped” directions of flow and spin, and the spin relaxation. These are parameterized by the spin Hall angle $\theta$, the swapping coefficient $\varkappa$, and the spin relaxation rate $\Gamma$, respectively. The external electromagnetic potential is introduced through the replacement $\partial_{\mu}\mapsto\partial_{\mu}-i[A_{\mu}\tau_{3},\cdot]$, and the Zeeman or exchange field is denoted by $\mathbf{h}$.
In equilibrium, this action can be related to the free energy $F[g]$ considered as a functional of the quasiclassical Green’s function [37]. At a given temperature $T$, the mapping $TS[Q]\mapsto iF[g]$ is achieved by replacing $\hat{\omega}_{tt^{\prime}}\rightarrow\omega_{n}$, the Matsubara frequency, and $Q\rightarrow g$. This connection becomes especially useful for superconductors as the minimization of $F[g]$ gives the equilibrium Usadel equation and determines the thermodynamics of the system.
@@
The first general symmetry of the NLSM is related to the redundancy of the Hilbert space in the Nambu description of superconductors, which involves both the full set of electronic states and their time-reversal conjugated counterparts. In nonsuperconducting systems, this technical doubling of the number of degrees of freedom is a convenient way to include fluctuations in the Cooperon channel. Because of the above redundancy, not all components of the matrix $Q$-field in NLSM are independent, but are constrained to satisfy the charge conjugation symmetry [26, 31, 32]. In the Keldysh contour representation, this fundamental symmetry condition reads $Q=C{Q}^{T}C^{-1}$, where $C=-{i}\sigma_{2}\tau_{1}$ is the charge conjugation operator, and the transposition operation involves both the usual matrix transposition and the interchanging of the contour time arguments  [^2] It is worth noting that the charge conjugation is an exact and universal symmetry of the Bogolyubov-de Gennes equations, which is the consequence of the redundancy of the Nambu formalism [70, 71, 72].. In the Keldysh matrix representation, after the standard Keldysh rotation [32] the charge conjugation symmetry of the 8$\times$8 matrix $Q$-field takes the form (see SM Sec. A for details),
@@
To illustrate this procedure, we derive contributions to the NLSM which may appear in the presence of SOC. More details are presented in SM Sec. C. Physically, SOC mixes the spin and orbital/translation degrees of freedom, resulting in the simultaneous presence of $\sigma$-matrices and spatial gradients in the NLSM. We construct the nontrivial contributions with the lowest orders on both Pauli matrices and gradients and discuss their main physical effects. Since SOC preserves the time-reversal, we only consider terms invariant under the operation $Q\mapsto\tau_{3}\sigma_{y}Q^{T}\sigma_{y}\tau_{3}$. Moreover, we consider only the dominant local in time contribution. Incorporation of small time-nonlocal terms breaking the particle-hole symmetry in NLSM is discussed in [45, 46].
Second order – spin precession and relaxation. In the second order, there are only three primitives consistent with the condition $Q^{2}=1$ and Eq. (2), namely, ${\rm Tr}(D_{kj}\partial_{k}Q\partial_{j}Q)$, ${\rm Tr}(\Gamma_{kj}\sigma_{k}Q\sigma_{j}Q)$, and ${\rm Tr}(\tilde{\alpha}_{kj}\sigma_{k}Q\partial_{j}Q)$. The chronological symmetry of Eq. (3) requires that $D_{kj}$ and $\Gamma_{kj}$ are real symmetric tensors, while $\tilde{\alpha}_{kj}=i\alpha_{kj}$ is an imaginary pseudotensor. $D_{kj}$ and $\Gamma_{kj}$ correspond, respectively, to the diffusion coefficient and the spin relaxation rate, which can be anisotropic, depending on the crystal symmetry. A second rank pseudotensor $\alpha_{kj}$ is allowed only in gyrotropic materials [47] and the corresponding term is responsible for the spin precession induced by SOC. By parameterizing $\alpha_{jk}$ as follows, $\alpha_{jk}=\mathcal{A}_{jl}D_{lk}$ the second order part of the NLSM can be cast in a compact form of SU(2) covariant diffusion with an additional spin relaxation,
$iS_{2}=\frac{\pi\nu}{8}{\rm Tr}\Big{(}-D_{kj}\tilde{\nabla}_{k}Q\tilde{\nabla}_{j}Q+\frac{1}{4}\tilde{\Gamma}_{kj}\sigma_{k}Q\sigma_{j}Q\Big{)}\;,$ (4)
where ~$\tilde{\nabla}_{k}=\partial_{k}-i[\hat{\mathcal{A}}_{k},\cdot]$ is the SU(2) covariant derivative with an effective gauge field $\hat{\mathcal{A}}_{k}=\mathcal{A}_{jk}\sigma_{j}$, and $\tilde{\Gamma}_{kj}=\Gamma_{kj}-4D_{ab}\mathcal{A}_{ka}\mathcal{A}_{jb}$  [^4] We note that stability of the system require $D_{kj}$ and $\Gamma_{kj}$ to be positive definite.. The SU(2) covariant diffusion has been introduced microscopically [11, 49] for a special class of $\mathbf{k}\cdot\mathbf{p}$ models with a Rashba-like linear SOC. Now we see that it is generic for gyrotropic materials. In realistic systems, the SU(2) symmetry of simple models is of course broken. However, at the leading order this leads only to extra spin relaxation, which can originate from the Dyakonov-Perel [50], or Elliott-Yafet [51, 52] mechanisms, or from the interplay of SOC with electron-electron interactions [53, 54].
Third order – spin-charge coupling related terms. There are two types of third order terms:
@@
$iS_{3}^{(2)}=\frac{\pi\nu}{16}{\rm Tr}\Big{(}-i\varepsilon_{ijl}\gamma_{lk}\sigma_{i}Q\sigma_{j}Q\partial_{k}Q+\beta_{ijk}\sigma_{i}Q\sigma_{j}\partial_{k}Q\Big{)}\;,$ (6)
@@
$\partial_{k}\mathcal{J}_{k}+[\tau_{3}(\hat{\omega}+i\mathbf{h}{\bm{\sigma}})+\hat{\Delta},g]=\mathcal{T}-\frac{1}{8}\Gamma_{jk}[\sigma_{j}g\sigma_{k},g]\;,$ (7)
@@
$\mathcal{J}_{k}$ $=-Dg\tilde{\nabla}_{k}g+\frac{i}{16}\varepsilon_{ijl}\gamma_{lk}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}\,,$ (8)
$\mathcal{T}$ $=i\alpha_{kj}[\sigma_{k},g\partial_{j}g]-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\,.$ (9)
By construction, these equations are consistent with the normalization condition $g^{2}=\mathbf{1}$ , where $\mathbf{1}$ is the identity matrix in Keldysh-Nambu-spin space. The retarded and advanced components of the general Usadel equation, Eq. (S93), reveal the spectral properties of the system, including spectral supercurrents. The Keldysh component provides a kinetic equation for distribution functions when the system is driven out of equilibrium. In the normal state, only the Keldysh component remains nontrivial 555 In Supplemental Material Sec. IV we derive from Eq.(S93) the charge and spin diffusion equations, which reveals explicitly all possible magnetoelectric effects allowed by symmetry.. It can be expressed through the distribution functions or through the physical observables such as charge and spin accumulation, see Eqs. (S71-S78) in SM Sec. D. Importantly, anticommutators of $g$ with spin matrices in Eqs. (S94)-(S95) induce coupling between singlet and triplet components of $g$. This indicates that the pseudotensor $\gamma_{ij}$ is related to magnetoelectric phenomena. The precise physics encoded in $\gamma_{ij}$ can be understood by analyzing Eqs. (S94)-(S95) in the linear regime.
@@
The second term in Eq. (10) describes a spin independent contribution to the matrix current proportional to the triplet (spin dependent) part of $\delta g$. Apparently, this corresponds to the spin-galvanic effect (SGE), also called the inverse Edelstein effect [2, 56, 57, 58, 59, 60, 61, 62, 63]. Indeed, the physical current is obtained from the Keldysh component of the matrix current $j_{k}=\frac{\pi\nu}{2}{\rm Tr}\{\tau_{3}\mathcal{J}_{k}^{K}(t,t)\}$ [64], which gives a linear relation between the charge current and the excess spin, which is defined as $\delta\mathbf{S}=\frac{\pi\nu}{2}{\rm Tr}\{\tau_{3}\mathbf{g}_{t}^{K}(t,t)\}$[64],
$j_{k}^{\rm SGE}=\gamma_{ik}\delta S_{i}\;.$ (12)
@@
In a vortex, the phase of the pair amplitude only depends on the angle, while their magnitude depends on the radial coordinate $r$. As elaborated in SM Sec. E, in Eq. (11) the phase derivative enters as an effective exchange field $\gamma_{kj}\hat{\phi}_{k}[\sigma_{j}\tau_{3},g_{s}]$, where $\bm{\hat{\phi}}$ denotes the unit vector in the tangential direction, while the derivative with respect to the radial coordinate does not break time-reversal symmetry and hence does not contribute to the spin texture. Thus, the spin around the vortex points along $\gamma_{kj}\hat{\phi}_{k}$. The inset of Fig. 1 sketches the textures for different point group symmetries. The dependence of the current density and induced spin density as a function of the radial coordinate, has to be computed numerically by solving the Usadel equation, see SM Sec. E for details. Fig. 1 shows the result for temperatures well below the critical temperature. The relation between the two is nonlocal, since $S(r)$ is maximized further away from the core than the maximum current density. This nonlocality is reduced by increasing the temperature.
@@
In the limit of high temperature, $T\lesssim T_{c}$ the GL functional can be straightforwardly derived [66, 67, 68, 9]. In a gyrotropic system, the term $q_{j}\gamma_{ij}$ acts as an effective exchange field, which combines with $\mathbf{h}_{i}$ (where $q_{k}$ represents the Fourier transform of the derivatives $i\partial_{k}$; see SM F). Consequently, the GL free energy includes a term linear in derivatives known as the Lifshitz invariant, denoted as $F_{L}=\chi_{ij}h_{j}\gamma_{ik}\partial_{k}\varphi$, where $\chi_{ij}$ is the magnetic response function to a Zeeman field, inducing an excess spin, $\delta S_{i}=\chi_{ij}h_{j}$. This establishes a connection of the Lifshitz invariant to the universal spin-galvanic coefficient $\gamma_{ij}$ appearing in Eq. (12) and in the action Eq. (6). The full list of all Lifshitz invariants for all gyrotropic point groups has been presented, for example, in Ref. [69].
In conclusion, we have presented a method for deriving the action of the Keldysh NLSM exclusively relying on the principles of charge conjugation, chronological and crystal symmetries. Applying this method, we have successfully derived the main objective of this paper, the quantum kinetic equation, commonly known as the Usadel equation, for metallic systems featuring generic SOC. Through this derivation, we have identified the distinct terms and symmetries governing spin precession, relaxation, and spin-charge coupling. As a byproduct, the provided NLSM can be used to investigate superconducting fluctuations or localization effects in materials with arbitrary types of SOC by applying various techniques developed in the context of effective field theories to address properties of NLSM beyond the saddle point level.
Our approach, while akin in spirit to the GL theory, offers an effective field theory valid in the whole temperature range. Moreover, it establishes a novel framework for deriving quantum kinetic theories applicable to diverse systems.